OpenStudy (muscrat123):

MEDAL AND FAN!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! i need some help. i learned it in class and i vaguely remember, but i cant remember much. ill post question in comments below. the directions are "identify each expression as a monomial, binomial, trinomial, or polynomial. write the degree of each expression." i know how to do this stuff really well, just not the fraction

OpenStudy (muscrat123):

\[9q+\frac{ 4q }{ 5p } - 3p ^{2}\]

OpenStudy (muscrat123):

@heretohelpalways

OpenStudy (muscrat123):

the directions are in the main comment'

OpenStudy (muscrat123):

question*

OpenStudy (anonymous):

@calculusxy

OpenStudy (anonymous):

@TTP

OpenStudy (muscrat123):

oh ok

OpenStudy (rizags):

ok, how many different terms do you see?

OpenStudy (rizags):

*let me rephrase that, How many different NONCOMBINABLE terms do you see?

OpenStudy (muscrat123):

i know it is a trinomial. and i know how to do the degree. i just dont know the fraction part

OpenStudy (rizags):

what exactly do you want to do to the fraction?

OpenStudy (rizags):

Its existence does not affect the answer

OpenStudy (muscrat123):

identify each expression as a monomial, binomial, trinomial, or polynomial. write the degree of each expression."

OpenStudy (rizags):

you mean each individual part of the expression?

OpenStudy (muscrat123):

i think. hang on. let me post the other ones i did

OpenStudy (muscrat123):

2z^2 - 5z - 10 = 2nd degree trinomial 24x^2y = 3rd degree monomial \[5a ^{2}b ^{3} - 4a ^{2}b ^{3}+2a ^{4}b ^{3}\] = 7th degree trinomial

OpenStudy (muscrat123):

@Mehek14

OpenStudy (anonymous):

I would say that the degree would be 2 for this trinomial. The first term has a degree of 1 given by q. The second term has degree 1/1, or still 1. The third term has a degree of 2. So the highest degree of the three terms is 2.

OpenStudy (anonymous):

The second term has a degree 0, actually. But that still does not affect our answer.

OpenStudy (anonymous):

If you are still confused, I would suggest visiting this page https://www.mathsisfun.com/algebra/degree-expression.html It's pretty helpful in explaining this topic

OpenStudy (muscrat123):

k thanks